# TabPFN: One Model to Rule Them All?

### v1

@@ Appendix C Covariate shift
Appendix D How does TabPFN extrapolate?
We evaluate the interpolation and extrapolation behavior of TabPFN trained on data sampled from predefined functions, comparing its performance to Gaussian process regression (GPR) as a baseline. Experiments are conducted in both one-dimensional (1D) and two-dimensional (2D) settings.
@@ D.1 Training data generation
1D case. The training set of size $n$ consists of features $x_{i}$ sampled uniformly over $[-1,1]$. Responses are generated as:
@@
    A randomly generated piecewise-linear function with $4$ pieces on $[-1,1]$.
TabPFN is trained on this small dataset and prediction is made over the extended interval $[-4,4]$ to assess both interpolation ($x\in[-1,1]$) and extrapolation ($x\notin[-1,1]$) performance.
2D case. The training set consists of features $\{(x_{i1},x_{i2}):i=1,\ldots,n\}$, where $x_{i1}$ and $x_{i2}$ are each sampled uniformly from $[-1,1]$. These points form a mesh grid, resulting in $n^{2}$ total observations. We examine two cases: $n=11$ and $n=31$. The responses are generated as:
@@
    Piecewise-bilinear function. Given a 2D grid of randomly generated function values $\{f_{ij}\}$ of size $4^{2}$. Uniformly divide $[-1,1]^{2}$ into $4^{2}$ cells. For $(x_{1},x_{2})$, find which cell it falls into. Suppose it falls into the $(i,j)$th cell, compute the local coordinates:
@@
TabPFN is trained on this small dataset and prediction is made over the uniform mesh grid over $[-4,4]^{2}$ to assess both interpolation ($x\in[-1,1]^{2}$) and extrapolation ($x\notin[-1,1]^{2}$) performance.
D.2 Gaussian process regression
For both 1D and 2D cases, we compare TabPFN with Gaussian process regression (GPR). For Gaussian process regression, we consider $5$ kernels: constant kernel $\times$ RBF kernel, constant kernel $\times$ Matern kernel, constant kernel $\times$ RationalQuadratic kernel, constant kernel $\times$ ExpSineSquared kernel, constant kernel $\times$ RBF kernel $+$ constant kernel $\times$ Matern kernel. The candidate variance $\sigma^{2}$ of the measurement error is over the grid $[0.05,0.1,0.15,0.2]$. We maximize the marginal log-likelihood to select the optimal kernel and its corresponding optimal parameters, and the optimal measurement error variance. The optimal parameter for GPR is implemented using the GaussianProcessRegressor in sklearn package in Python.
D.3 Results
@@
Appendix E Additional results for sparse linear regression
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: Accumulated Local Effects plot for TabPFN under sparse linear regression when $d=100$ under orthogonal design. Columns (left to right) correspond to increasing sparsity levels (1, 5, 10, 20, 30). Rows 1–4 show results when training set is of size $n=50$. First row: Relevant feature; second row: randomly sampled $5$ irrelevant features (SNR = 1.22). Third row: Relevant feature; fourth row: randomly sampled $5$ irrelevant features (SNR = 6). Rows 5–8 show the same results when training set is of size $n=500$.
@@ Appendix F Implementation Details for LDA and kNN
For $k$-nearest neighbors ($k$NN) classification, we select the optimal number of neighbors $k$ via $5$-fold cross-validation. The search grid consists of $10$ equally spaced integers between $\lfloor n^{1/4}\rfloor$ and $\lfloor n^{3/4}\rfloor$, with the value maximizing cross-validation accuracy chosen for each experimental replication.

### v2

@@ C.1 Additional simulation results
C.2 Real data comparison
Airfoil Self-Noise Dataset. The Airfoil dataset, used previously in studies on covariate shift [51], contains 1,503 observations with a five-dimensional covariate vector (log frequency, angle of attack, chord length, free-stream velocity, and suction-side log displacement thickness) and a continuous response variable $Y$. The data are randomly divided into training and testing subsets with a 7:3 ratio, that is, $70\%$ of the observations are used for training and $30\%$ for testing. To induce covariate shift, we follow the procedure in Tibshirani et al. [51]: we first split the data as described, and then resample the test set with replacement using importance weights
$\omega(\bm{x},y)=\exp(\bm{x}^{\mathrm{T}}\bm{\beta}).$
We consider two different shift settings characterized by
$\bm{\beta}_{1}=[-1,1,-1,1,-1]^{\mathrm{T}},\quad\bm{\beta}_{2}=[-1,0,0,0,1]^{\mathrm{T}}.$
For each $\bm{\beta}$ setting, the data splitting and resampling procedure is repeated ten times, and the mean squared prediction error (MSE) is averaged over the ten repetitions.
Concrete Compressive Strength Dataset. The Concrete dataset consists of 1,030 observations with an eight-dimensional covariate vector (cement, blast furnace slag, fly ash, water, superplasticizer, coarse aggregate, fine aggregate, and age) and the target variable is the measured compressive strength of concrete. As with the Airfoil data, the dataset is randomly split into $70\%$ training and $30\%$ testing observations. Covariate shift is then introduced by resampling the test set with replacement using
$\omega(\bm{x},y)=\exp(\bm{x}^{\mathrm{T}}\bm{\beta}),$
and we again consider two distinct settings:
$\bm{\beta}_{1}=[-1,1,-1,1,-1,1,-1,1]^{\mathrm{T}},\quad\bm{\beta}_{2}=[-1,0,0,0,0,0,0,1]^{\mathrm{T}}.$
Each $\bm{\beta}$ setting is repeated for ten random data splits, and the average MSE is reported.
Methods and Evaluation. We compare the proposed TabPFN approach with several baseline methods (same as in the simulation study), including the naive gradient boosting regression, its importance-weighted variant, and the kernel ridge regression (KRR) methods from Wang [54]. Since the code in Wang [54] was limited to univariate $X$, we extend it to the multivariate case using a radial basis function (RBF) kernel. For each method and setting, the reported $\pm$ values denote the standard error of the prediction mean squared error (PMSE) computed across the ten repetitions.
Appendix D Comparison with Gaussian process regression
To study the adaptiveness of TabPFN, we evaluate the interpolation and extrapolation behavior of TabPFN trained on data sampled from predefined functions, comparing its performance to Gaussian process regression (GPR) as a baseline. Experiments are conducted in both one-dimensional (1D) and two-dimensional (2D) settings.
@@ D.1 Training data generation
1D case. The training set of size $n$ consists of features $x_{i}$ sampled uniformly over [1,1][-1,1]. Responses are generated as:
@@
    A randomly generated piecewise-linear function with $4$ pieces on [1,1][-1,1].
TabPFN is trained on this small dataset and prediction is made over the extended interval [4,4][-4,4] to assess both interpolation ($x\in[-1,1]$) and extrapolation ($x\notin[-1,1]$) performance.
2D case. The training set consists of features $\{(x_{i1},x_{i2}):i=1,\ldots,n\}$, where $x_{i1}$ and $x_{i2}$ are each sampled uniformly from [1,1][-1,1]. These points form a mesh grid, resulting in $n^{2}$ total observations. We examine two cases: $n=11$ and $n=31$. The responses are generated as:
@@
    Piecewise-bilinear function. Given a 2D grid of randomly generated function values $\{f_{ij}\}$ of size $4^{2}$. Uniformly divide $1,1]^{2}$ into $4^{2}$ cells. For $(x_{1},x_{2})$, find which cell it falls into. Suppose it falls into the $(i,j)$th cell, compute the local coordinates:
@@
TabPFN is trained on this small dataset and prediction is made over the uniform mesh grid over $4,4]^{2}$ to assess both interpolation ($x\in[-1,1]^{2}$) and extrapolation ($x\notin[-1,1]^{2}$) performance.
D.2 Baseline for comparison: Gaussian process regression
For both 1D and 2D cases, we compare TabPFN with Gaussian process regression (GPR). For Gaussian process regression, we consider $5$ kernels: constant kernel $\times$ RBF kernel, constant kernel $\times$ Matern kernel, constant kernel $\times$ RationalQuadratic kernel, constant kernel $\times$ ExpSineSquared kernel, constant kernel $\times$ RBF kernel ++ constant kernel $\times$ Matern kernel. The candidate variance $\sigma^{2}$ of the measurement error is over the grid $[0.05,0.1,0.15,0.2]$. We maximize the marginal log-likelihood to select the optimal kernel and its corresponding optimal parameters, and the optimal measurement error variance. The GPR is implemented using the GaussianProcessRegressor in scikit-learn package in Python.
D.3 Interpolation and extrapolation
The results for $n=31$ training data points are shown in Figure 16.
@@ D.3 Interpolation and extrapolation
We observe that in all cases, TabPFN fits a smoothly-varying curve that broadly resembles the GPR curve within the interpolation region. The notable differences are:
@@
D.4 Uncertainty quantification
To assess the uncertainty quantification capability of TabPFN in comparison with GPR, we evaluate the predictive intervals of both methods in terms of their coverage ratio and interval length. For each predefined function, we generate $100$ independent datasets following the data generation procedure described above. For each dataset, the model is trained on the corresponding training set and evaluated on a dense grid with a step size of $0.1$ over the range [4.5,4.6][-4.5,4.6], which encompasses both the interpolation region ([1,1][-1,1]) and the extrapolation region ($4.5,4.6]\setminus[-1,1]$). At each grid point, we record the predictive mean and the associated $95\%$ prediction interval produced by the model. For TabPFN, the $95\%$ prediction interval is obtained from the empirical $2.5\%$ and $97.5\%$ quantiles of the posterior predictive distribution. The coverage ratio is computed as the proportion of test points whose true function values fall within the predicted intervals, averaged over all repetitions. Similarly, the interval length is computed as the mean width of the predictive intervals, also averaged across all test points and repetitions. This evaluation jointly examines the calibration (via coverage ratio) and sharpness (via interval length) of the uncertainty estimates provided by TabPFN and GPR in both interpolation and extrapolation settings. The average coverage ratios and interval lengths for interpolation and extrapolation are reported in Table 9 and Table 10, respectively.
Table 9: Average coverage ratio and interval length of TabPFN and Gaussian process regression (GPR) under the interpolation setting ($x\in[-1,1]$), averaged over 100 repetitions.
$n$ Method Linear Quadratic Step P Linear
Coverage Length Coverage Length Coverage Length Coverage Length
11 TabPFN 0.9350 0.4656 0.9525 0.7365 0.9705 1.2301 0.9930 1.8449
GPR 0.9455 0.4304 0.9665 0.5439 0.8635 0.6934 0.9795 0.6977
31 TabPFN 0.9405 0.3924 0.9420 0.4235 0.9365 0.5501 0.9375 0.4788
GPR 0.8775 0.3268 0.9390 0.4247 0.9050 0.5808 0.9550 0.5268
Table 10: Average coverage ratio and interval length of TabPFN and Gaussian process regression (GPR) under the extrapolation setting ($x\notin[-1,1]$), averaged over 100 repetitions.
$n$ Method Linear Quadratic Step P Linear
Coverage Length Coverage Length Coverage Length Coverage Length
11 TabPFN 0.9276 5.1478 0.4287 8.4742 0.9996 3.3699 1.0000 2.4744
GPR 0.4856 2.9333 0.0792 2.7310 0.9997 2.4065 1.0000 1.7769
31 TabPFN 0.9827 5.9242 0.6817 11.9492 0.9975 3.8227 0.9993 4.9580
GPR 0.6175 2.6130 0.0882 3.2259 0.9999 2.4096 1.0000 1.7710
Interpolation performance (Table 9). In the interpolation region ($x\in[-1,1]$), TabPFN consistently exhibits coverage at around the nominal 95% level across all function types, indicating that it remains well-calibrated within the training distribution. In contrast, GPR fails to achieve the desired coverage under the step function, suggesting difficulty in handling discontinuities. Regarding interval length, TabPFN generally produces wider intervals than GPR when $n=11$, particularly for nonlinear functions such as the quadratic and piecewise-linear cases, implying more conservative uncertainty estimates with limited data. As the training sample size increases to $n=31$, TabPFN’s intervals become narrower and comparable to those of GPR, reflecting sharper and still well-calibrated uncertainty estimates as data availability improves.
Extrapolation performance(Table 10). In contrast, the extrapolation region ($x\notin[-1,1]$) reveals a clear divergence between the two models. TabPFN maintains substantially higher coverage than GPR in nearly all cases. For smooth functions such as the linear and quadratic cases, GPR’s coverage drops sharply (e.g., below 0.5 for the linear and below 0.1 for the quadratic function), indicating severe overconfidence and poor uncertainty calibration outside the training range. TabPFN, however, achieves much higher coverage (0.93–0.98 for linear; 0.43–0.68 for quadratic) and adapts its interval width dynamically to reflect function complexity—producing wider intervals in regions of greater uncertainty (e.g., the quadratic case). For discontinuous or piecewise functions (step and piecewise-linear), both models attain full coverage, though TabPFN generally yields slightly longer intervals, implying a more conservative but robust uncertainty quantification. Overall, TabPFN exhibits markedly better-calibrated uncertainty estimates than GPR under extrapolation, effectively avoiding the overconfident predictions that GPR commonly produces outside its training support.
Summary. These results demonstrate that while TabPFN and GPR perform comparably under interpolation, TabPFN provides superior extrapolative uncertainty quantification, maintaining both coverage and adaptive interval width even in regions far from the training data.
Appendix E Sparse linear regression
E.1 Hyper-parameter selection
The LASSO and Ridge regression models are implemented using the Lasso and Ridge functions in the scikit-learn package, while the MCP and SCAD estimators are implemented using the pycasso package in Python. All four methods involve a regularization hyperparameter controlling the strength of penalization. For each method, we evaluate $20$ candidate values of the tuning parameter and select the optimal one based on the prediction error computed on an independent validation set of size $1000$. For LASSO, the regularization grid is chosen over the interval $0.001\|X_{\text{train}}^{\top}y_{\text{train}}\|_{\infty},\|X_{\text{train}}^{\top}y_{\text{train}}\|_{\infty}]$. For Ridge, SCAD, and MCP, the tuning parameter grids are respectively defined as $[0.01,100]$, $[3,100]$, and $[1.1,100]$. All grids are generated on a logarithmic scale with $20$ evenly spaced points within the corresponding intervals.
E.2 Additional results for sparse linear regression
The results under non-orthogonal design with beta-type 1 and beta-type 2 are given respectively in Figure 19 and Figure 20.
Figure 19: Relative test MSE of various methods compared to TabPFN. Experiments use beta-type 1 with nonorthogonal design. The top and bottom rows display results for $n=50$ and $n=500$ respectively, while the columns (from left to right) correspond to increasing SNR values.
Figure 20: Relative test MSE of various methods compared to TabPFN. Experiments use beta-type 2 with nonorthogonal design. The top and bottom rows display results for $n=50$ and $n=500$ respectively, while the columns (from left to right) correspond to increasing SNR values.
E.3 ALE plot for TabPFN
@@
Figure 21: Accumulated Local Effects plot for TabPFN under sparse linear regression when $d=100$ under orthogonal design. Columns (left to right) correspond to increasing sparsity levels (1, 5, 10, 20, 30). Rows 1–4 show results when training set is of size $n=50$. First row: Relevant feature; second row: randomly sampled 55 irrelevant features (SNR = 1.22). Third row: Relevant feature; fourth row: randomly sampled 55 irrelevant features (SNR = 6). Rows 5–8 show the same results when training set is of size $n=500$.
@@ Appendix F Implementation Details for LDA and kNN
For $k$-nearest neighbors (kkNN) classification, we select the optimal number of neighbors $k$ via $5$-fold cross-validation. The search grid consists of $10$ equally spaced integers between $\lfloor n^{1/4}\rfloor$ and $\lfloor n^{3/4}\rfloor$, with the value maximizing cross-validation accuracy chosen for each experimental replication.
{funding}